Tony runs a cell phone kiosk in the mall and has 220 potential customers each day. He has a 6 % chance to make a sale to each customer. Use continuity correction to normally approximate the following probabilities. Round to 4 decimal places.

a. The probability that he makes AT LEAST 14 sales

b. The probability that he makes AT MOST 12 sales
Answer #1

a)

n= 220 p= 0.0600
here mean of distribution=μ=np= 13.2
and standard deviation σ=sqrt(np(1-p))= 3.5225

probability that he makes AT LEAST 14 sales

probability = P(X>13.5) = P(Z>0.09)= 1-P(Z<0.09)= 1-0.5359= 0.4641

b)

probability that he makes AT MOST 12 sales :

probability = P(X<12.5) = P(Z<-0.2)= 0.4207